# Deep Learning for 3D Human Pose Estimation and Mesh Recovery: A Survey

### v1

@@ 6.1 Evaluation metrics
Mean Per Joint Localization Error (MPJLE) is a more perceptive and robust evaluation metric than MPJPE and MPJAE ionescu_human36m_2014 . It allows for an adjustable tolerance level through a perceptual threshold $t$ :
@@
There are various modifications of MPJPE, MPJAE, and MPJLE, including Procrustes-aligned (PA-) MPJPE, MPJAE, MPJLE and Normalized (N-) MPJPE, MPJAE, MPJLE. The former refers to procrustes-aligned metrics, while the latter denotes normalized metrics. Additionally, some 2D metrics are adaptable to 3D, such as the 3D Percentage of Correct Keypoints (3D PCK) and the 3D Area Under the Curve (3D AUC). The PCK YangandRamanan2013 measures the distance between predicted and groundtruth keypoints, considering keypoints correct if this distance is less than a predefined threshold. The AUC signifies the aggregate area beneath the PCK threshold curve as the threshold varies.
@@ 6.2 Datasets
Human3.6M ionescu_human36m_2014 3D/Mesh Video 3.6M multi-view Website
3DPW von2018recovering 3D/Mesh Video 51K multi-person Website
MPI-INF-3DPH mehta2017monocular 2D/3D Video 2K in-wild Website
HumanEva sigal2010humaneva 3D Video 40K multi-view Website
CMU-Panoptic Joo_2015_ICCV 3D Video 1.5M multi-view/multi-person Website
MuCo-3DHP mehta2018single 3D Image 8K multi-person/occluded scence Website
SURREAL varol2017learning 2D/3D/Mesh Video 6.0M synthetic model Website
3DOH50K zhang2020object 2D/3D/Mesh Image 51K object-occluded Website
3DCP muller2021self Mesh Mesh 190 contact Website
AMASS mahmood2019amass Mesh Motion 11K soft-tissue dynamics Website
DensePose guler2018densepose Mesh Image 50K multi-person Website
UP-3D lassner2017unite 3D/Mesh Image 8K sport scence Website
THuman2.0 zheng2019deephuman Mesh Image 7K textured surface Website
@@
Human3.6M dataset ionescu_human36m_2014 is the most widely used dataset in the evaluation of 3D human pose estimation. It encompasses a vast collection of 3.6 million poses captured using RGB and ToF cameras from diverse viewpoints within a real-world setting. Additionally, this dataset incorporates high-resolution 3D scanner data of body meshes, playing a pivotal role in the progression of human sensing systems. Table 4 exhibits the performance of state-of-the-art 3D human pose estimation methods on the Human3.6M dataset.
Table 4: Comparisons of 3D pose estimation methods on Human3.6M ionescu_human36m_2014 .
Method Year Publication Highlight MPJPEnormal-$\downarrow$ PMPJPEnormal-$\downarrow$ Code
Graformer zhao2022graformer 2022 CVPR’22 graph-based transformer 35.2 - Code
GLA-GCN yu2023gla 2023 ICCV’23 adaptive GCN 34.4 37.8 Code
PoseDA chai2023global 2023 arXiv’23 domain adaptation 49.4 34.2 Code
GFPose ci2023gfpose 2023 CVPR’23 gradient fields 35.6 30.5 Code
TP-LSTMs lee2022human 2022 TPAMI’22 pose similarity metric 40.5 31.8 -
FTCM tang2023ftcm 2023 TCSVT’23 frequency-temporal collaborative 28.1 - Code
VideoPose3D pavllo20193d 2019 CVPR’19 semi-supervised 46.8 36.5 Code
PoseFormer zheng20213d 2021 ICCV’21 spatio-temporal transformer 44.3 34.6 Code
STCFormer tang20233d 2023 CVPR’23 spatio-temporal transformer 40.5 31.8 Code
3Dpose_ssl wang20193d 2020 TPAMI’20 self-supervised 63.6 63.7 Code
MTF-Transformer shuai2022adaptive 2022 TPAMI’22 multi-view temporal fusion 26.2 - Code
AdaptPose gholami2022adaptpose 2022 CVPR’22 cross datasets 42.5 34.0 Code
3D-HPE-PAA xue2022boosting 2022 TIP’22 part aware attention 43.1 33.7 Code
DeciWatch zeng2022deciwatch 2022 ECCV’22 efficient framework 52.8 - Code
Diffpose gong2023diffpose 2023 CVPR’23 pose refine 36.9 28.7 Code
Elepose wandt2022elepose 2022 CVPR’22 unsupervised - 36.7 Code
Uplift and Upsample einfalt2023uplift 2023 CVPR’23 efficient transformers 48.1 37.6 Code
RS-Net hassan2023regular 2023 TIP’23 regular splitting graph network 48.6 38.9 Code
HSTFormer qian2023hstformer 2023 arXiv’23 spatial-temporal transformers 42.7 33.7 Code
PoseFormerV2 zhao2023poseformerv2 2023 CVPR’23 frequency domain 45.2 35.6 Code
DiffPose holmquist2023diffpose 2023 ICCV’23 diffusion models 42.9 30.8 Code
@@
MPI-INF-3DPH dataset mehta2017monocular offers over 2K videos featuring joint annotations of 13 keypoints in outdoor scenes, apt for 2D and 3D human pose estimation. The ground truth was obtained using a multi-camera arrangement and a marker-less MoCap system, representing a shift from traditional marker-based MoCap systems that involve real individuals. Table 5 showcases the performance of state-of-the-art methods on the 3DPH dataset.
Table 5: Comparisons of 3D pose estimation methods on MPI-INF-3DPH mehta2017monocular .
Method Year Publication Highlight MPJPEnormal-$\downarrow$ PCKnormal-$\uparrow$ AUCnormal-$\uparrow$ Code
HSTFormer qian2023hstformer 2023 arXiv’23 spatial-temporal transformers 28.3 98.0 78.6 Code
PoseFormerV2 zhao2023poseformerv2 2023 CVPR’23 frequency domain 27.8 97.9 78.8 Code
Uplift and Upsample einfalt2023uplift 2023 CVPR’23 efficient transformers 46.9 95.4 67.6 Code
RS-Net hassan2023regular 2023 TIP’23 regular splitting graph network - 85.6 53.2 Code
Diffpose gong2023diffpose 2023 CVPR’23 pose refine 29.1 98.0 75.9 Code
FTCM tang2023ftcm 2023 TCSVT’23 frequency-temporal collaborative 31.2 97.9 79.8 Code
STCFormer tang20233d 2023 CVPR’23 spatio-temporal transformer 23.1 98.7 83.9 Code
PoseDA chai2023global 2023 arXiv’23 domain adaptation 61.3 92.0 62.5 Code
TP-LSTMs lee2022human 2022 TPAMI’22 pose similarity metric 48.8 82.6 81.3 -
AdaptPose gholami2022adaptpose 2022 CVPR’22 cross datasets 77.2 88.4 54.2 Code
3D-HPE-PAA xue2022boosting 2022 TIP’22 part aware attention 69.4 90.3 57.8 Code
Elepose wandt2022elepose 2022 CVPR’22 unsupervised 54.0 86.0 50.1 Code
@@
3DPW dataset von2018recovering captures 51,000 sequences of single-view video, complemented by IMUs data. These videos were recorded using a handheld camera, with the IMU data facilitating the association of 2D poses with their 3D counterparts. 3DPW stands out as one of the most formidable datasets, establishing itself as a benchmark for 3D pose estimation in multi-person, in-wild scenarios of recent times. Table 6 shows the performance of state-of-the-art human mesh recovery methods on the Human3.6M and 3DPW datasets.
Table 6: Comparisons of human mesh recovery methods on Human3.6M ionescu_human36m_2014 and 3DPW von2018recovering .
@@
MPJPEnormal-$\downarrow$ PA-MPJPEnormal-$\downarrow$ MPJPEnormal-$\downarrow$ PA-MPJPEnormal-$\downarrow$ PVEnormal-$\downarrow$
VirtualMarker ma20233d CVPR’23 novel intermediate representation 47.3 32.0 67.5 41.3 77.9 Code
NIKI li2023niki CVPR’23 inverse kinematics - - 71.3 40.6 86.6 Code
TORE dou2023tore ICCV’23 efficient transformer 59.6 36.4 72.3 44.4 88.2 Code
JOTR li2023jotr ICCV’23 contrastive learning - - 76.4 48.7 92.6 Code
HMDiff foo2023distribution ICCV’23 reverse diffusion processing 49.3 32.4 72.7 44.5 82.4 Code
ReFit wang2023refit ICCV’23 recurrent fitting network 48.4 32.2 65.8 41.0 - Code
@@
PyMAF-X zhang2023pymaf TPAMI’23 regression-based one-stage whole body - - 74.2 45.3 87.0 Code
PointHMR kim2023sampling CVPR’23 vertex-relevant feature extraction 48.3 32.9 73.9 44.9 85.5 -
PLIKS shetty2023pliks CVPR’23 inverse kinematics 47.0 34.5 60.5 38.5 73.3 Code
ProPose fang2023learning CVPR’23 learning analytical posterior probability 45.7 29.1 68.3 40.6 79.4 Code
POTTER zheng2023potter CVPR’23 pooling attention transformer 56.5 35.1 75.0 44.8 87.4 Code
PoseExaminer liu2023poseexaminer ICCV’23 automated testing of out-of-distribution - - 74.5 46.5 88.6 Code
MotionBERT zhu2023motionbert ICCV’23 pretrained human representations 43.1 27.8 68.8 40.6 79.4 Code
3DNBF zhang20233d ICCV’23 analysis-by-synthesis approach - - 88.8 53.3 - Code
FastMETRO cho2022cross ECCV’22 efficient architecture 52.2 33.7 73.5 44.6 84.1 Code
CLIFF li2022cliff ECCV’22 multi-modality inputs 47.1 32.7 69.0 43.0 81.2 Code
PARE kocabas2021pare ICCV’21 part-driven attention - - 74.5 46.5 88.6 Code
Graphormer lin2021mesh ICCV’21 GCNN-reinforced transformer 51.2 34.5 74.7 45.6 87.7 Code
PSVT qiu2023psvt CVPR’23 spatio-temporal encoder - - 73.1 43.5 84.0 -
@@
GLoT shen2023global CVPR’23 short-term and long-term temporal correlations 67.0 46.3 80.7 50.6 96.3 Code
MPS-Net wei2022capturing CVPR’23 temporally adjacent representations 69.4 47.4 91.6 54.0 109.6 Code
MAED wan2021encoder ICCV’21 multi-level attention 56.4 38.7 79.1 45.7 92.6 Code
Lee et al. lee2021uncertainty ICCV’21 uncertainty-aware 58.4 38.4 92.8 52.2 106.1 -
TCMR choi2021beyond CVPR’21 temporal consistency 62.3 41.1 95.0 55.8 111.3 -
VIBE kocabas2020vibe CVPR’20 self-attention temporal network 65.6 41.4 82.9 51.9 99.1 Code
ImpHMR cho2023implicit CVPR’23 implicitly imagine person in 3D space - - 74.3 45.4 87.1 -
SGRE wang20233d ICCV’23 sequentially global rotation estimation - - 78.4 49.6 93.3 Code
PMCE you2023co ICCV’23 pose and mesh co-evolution network 53.5 37.7 69.5 46.7 84.8 Code
@@
HumanEva dataset sigal2010humaneva is a multi-view 3D human pose estimation dataset comprising two versions: HumanEva-I and HumanEva-II. In HumanEva-I, the dataset includes around 40,000 multi-view video frames captured from seven cameras positioned at the front, left, and right (RGB) and four corners (Mono). Additionally, HumanEva-II features approximately 2,460 frames, recorded with four cameras at each corner.
@@
CMU-Panoptic dataset Joo_2015_ICCV ; simon2017hand includes 65 frame sequences, approximately 5.5 hours of footage, and features 1.5 million 3D annotated poses. Recorded via a massively multi-view system equipped with 511 calibrated cameras and 10 RGB-D sensors featuring hardware-based synchronization, this dataset is crucial for developing weakly supervised methods through multi-view geometry. These methods address the occlusion problems commonly encountered in traditional computer vision techniques.
@@
Multiperson Composited 3D Human Pose (MuCo-3DHP) dataset mehta2018single serves as a large-scale, multi-person occluded training set for 3D human pose estimation. Frames in the MuCo-3DHP are generated from the MPI-INF-3DPH dataset through a compositing and augmentation scheme.
@@
SURREAL dataset varol2017learning is a large synthetic human body dataset containing 6 million RGB video frames. It provides a range of accurate annotations, including depth, body parts, optical flow, 2D/3D poses, and surfaces. In the SURREAL dataset, images exhibit variations in texture, view, and pose, and the body models are based on the SMPL parameters, a widely-recognized mesh representation standard.
@@
3DOH50K dataset zhang2020object offers a collection of 51,600 images obtained from six distinct viewpoints in real-world settings, predominantly featuring object occlusions. Each image is annotated with ground truth 2D and 3D poses, SMPL parameters, and a segmentation mask. Utilized for training human estimation and reconstruction models, the 3DOH50K dataset facilitates exceptional performance in occlusion scenarios.
@@
3DCP dataset muller2021self represents a 3D human mesh dataset, derived from AMASS mahmood2019amass . It includes 190 self-contact meshes spanning six human subjects (three males and three females), each modeled with an SMPL-X parameterized template.
@@
AMASS dataset mahmood2019amass constitutes a comprehensive and diverse human motion dataset, encompassing over 11,000 motions from 300 subjects, totaling more than 40 hours. The motion data, accompanied by SMPL parameters for skeleton and mesh representation, is derived from a marker-based MoCap system utilizing 15 optical markers.
@@
DensePose dataset guler2018densepose features 50,000 manually annotated real images, comprising 5 million image-to-surface correspondence pairs extracted from the COCO lin2014microsoft dataset. This dataset proves instrumental for training in dense human pose estimation, as well as in detection and segmentation tasks.
@@
UP-3D dataset lassner2017unite is a dedicated 3D human pose and shape estimation dataset featuring extensive annotations in sports scenarios. The UP-3D comprises approximately 8,000 images from the LSP and MPII datasets. Additionally, each image in UP-3D is accompanied by a metadata file indicating the quality (medium or high) of the 3D fit.
@@
THuman dataset zheng2019deephuman constitutes a 3D real-world human mesh dataset. It includes 7,000 RGBD images, each featuring a textured surface mesh obtained using a Kinect camera. Including surface mesh with detailed texture and the aligned SMPL model is anticipated to significantly enhance and stimulate future research in human mesh reconstruction.

### v2

@@ 6.1 Evaluation metrics
Mean Per Joint Localization Error (MPJLE) is a more perceptive and robust evaluation metric than MPJPE and MPJAE [225]. It allows for an adjustable tolerance level through a perceptual threshold $t$ :
@@
There are various modifications of MPJPE, MPJAE, and MPJLE, including Procrustes-aligned (PA-) MPJPE, MPJAE, MPJLE and Normalized (N-) MPJPE, MPJAE, MPJLE. The former refers to procrustes-aligned metrics, while the latter denotes normalized metrics. Additionally, some 2D metrics are adaptable to 3D, such as the 3D Percentage of Correct Keypoints (3D PCK) and the 3D Area Under the Curve (3D AUC). The PCK [226] measures the distance between predicted and groundtruth keypoints, considering keypoints correct if this distance is less than a predefined threshold. The AUC signifies the aggregate area beneath the PCK threshold curve as the threshold varies.
@@ 6.2 Datasets
Human3.6M [225] 3D/Mesh Video 3.6M multi-view Website
3DPW [227] 3D/Mesh Video 51K multi-person Website
MPI-INF-3DHP [228] 2D/3D Video 2K in-wild Website
HumanEva [229] 3D Video 40K multi-view Website
CMU-Panoptic [230] 3D Video 1.5M multi-view/multi-person Website
MuCo-3DHP [118] 3D Image 8K multi-person/occluded scence Website
SURREAL [231] 2D/3D/Mesh Video 6.0M synthetic model Website
3DOH50K [153] 2D/3D/Mesh Image 51K object-occluded Website
3DCP [232] Mesh Mesh 190 contact Website
AMASS [233] Mesh Motion 11K soft-tissue dynamics Website
DensePose [234] Mesh Image 50K multi-person Website
UP-3D [235] 3D/Mesh Image 8K sport scence Website
THuman2.0 [236] Mesh Image 7K textured surface Website
@@
Human3.6M dataset [225] is the most widely used dataset in the evaluation of 3D human pose estimation. It encompasses a vast collection of 3.6 million poses captured using RGB and ToF cameras from diverse viewpoints within a real-world setting. Additionally, this dataset incorporates high-resolution 3D scanner data of body meshes, playing a pivotal role in the progression of human sensing systems. Table 4 exhibits the performance of state-of-the-art 3D human pose estimation methods on the Human3.6M dataset.
Table 4: Comparisons of 3D pose estimation methods on Human3.6M [225].
[…]
Graformer [237] 2022 CVPR’22 graph-based transformer 35.2 - Code
GLA-GCN [238] 2023 ICCV’23 adaptive GCN 34.4 37.8 Code
PoseDA [49] 2023 arXiv’23 domain adaptation 49.4 34.2 Code
GFPose [239] 2023 CVPR’23 gradient fields 35.6 30.5 Code
TP-LSTMs [240] 2022 TPAMI’22 pose similarity metric 40.5 31.8 -
FTCM [122] 2023 TCSVT’23 frequency-temporal collaborative 28.1 - Code
VideoPose3D [89] 2019 CVPR’19 semi-supervised 46.8 36.5 Code
PoseFormer [90] 2021 ICCV’21 spatio-temporal transformer 44.3 34.6 Code
STCFormer [96] 2023 CVPR’23 spatio-temporal transformer 40.5 31.8 Code
3Dpose_ssl [85] 2020 TPAMI’20 self-supervised 63.6 63.7 Code
MTF-Transformer [32] 2022 TPAMI’22 multi-view temporal fusion 26.2 - Code
AdaptPose [88] 2022 CVPR’22 cross datasets 42.5 34.0 Code
3D-HPE-PAA [101] 2022 TIP’22 part aware attention 43.1 33.7 Code
DeciWatch [241] 2022 ECCV’22 efficient framework 52.8 - Code
Diffpose [242] 2023 CVPR’23 pose refine 36.9 28.7 Code
Elepose [83] 2022 CVPR’22 unsupervised - 36.7 Code
Uplift and Upsample [21] 2023 CVPR’23 efficient transformers 48.1 37.6 Code
RS-Net [59] 2023 TIP’23 regular splitting graph network 48.6 38.9 Code
HSTFormer [95] 2023 arXiv’23 spatial-temporal transformers 42.7 33.7 Code
PoseFormerV2 [60] 2023 CVPR’23 frequency domain 45.2 35.6 Code
DiffPose [243] 2023 ICCV’23 diffusion models 42.9 30.8 Code
@@
MPI-INF-3DHP dataset [228] offers over 2K videos featuring joint annotations of 13 keypoints in outdoor scenes, apt for 2D and 3D human pose estimation. The ground truth was obtained using a multi-camera arrangement and a marker-less MoCap system, representing a shift from traditional marker-based MoCap systems that involve real individuals. Table 5 showcases the performance of state-of-the-art methods on the 3DHP dataset.
Table 5: Comparisons of 3D pose estimation methods on MPI-INF-3DHP [228].
[…]
HSTFormer [95] 2023 arXiv’23 spatial-temporal transformers 28.3 98.0 78.6 Code
PoseFormerV2 [60] 2023 CVPR’23 frequency domain 27.8 97.9 78.8 Code
Uplift and Upsample [21] 2023 CVPR’23 efficient transformers 46.9 95.4 67.6 Code
RS-Net [59] 2023 TIP’23 regular splitting graph network - 85.6 53.2 Code
Diffpose [242] 2023 CVPR’23 pose refine 29.1 98.0 75.9 Code
FTCM [122] 2023 TCSVT’23 frequency-temporal collaborative 31.2 97.9 79.8 Code
STCFormer [96] 2023 CVPR’23 spatio-temporal transformer 23.1 98.7 83.9 Code
PoseDA [49] 2023 arXiv’23 domain adaptation 61.3 92.0 62.5 Code
TP-LSTMs [240] 2022 TPAMI’22 pose similarity metric 48.8 82.6 81.3 -
AdaptPose [88] 2022 CVPR’22 cross datasets 77.2 88.4 54.2 Code
3D-HPE-PAA [101] 2022 TIP’22 part aware attention 69.4 90.3 57.8 Code
Elepose [83] 2022 CVPR’22 unsupervised 54.0 86.0 50.1 Code
@@
3DPW dataset [227] captures 51,000 sequences of single-view video, complemented by IMUs data. These videos were recorded using a handheld camera, with the IMU data facilitating the association of 2D poses with their 3D counterparts. 3DPW stands out as one of the most formidable datasets, establishing itself as a benchmark for 3D pose estimation in multi-person, in-wild scenarios of recent times. Table 6 shows the performance of state-of-the-art human mesh recovery methods on the Human3.6M and 3DPW datasets.
Table 6: Comparisons of human mesh recovery methods on Human3.6M [225] and 3DPW [227].
@@
[…]
VirtualMarker [244] CVPR’23 novel intermediate representation 47.3 32.0 67.5 41.3 77.9 Code
NIKI [160] CVPR’23 inverse kinematics - - 71.3 40.6 86.6 Code
TORE [149] ICCV’23 efficient transformer 59.6 36.4 72.3 44.4 88.2 Code
JOTR [170] ICCV’23 contrastive learning - - 76.4 48.7 92.6 Code
HMDiff [202] ICCV’23 reverse diffusion processing 49.3 32.4 72.7 44.5 82.4 Code
ReFit [165] ICCV’23 recurrent fitting network 48.4 32.2 65.8 41.0 - Code
@@
PyMAF-X [184] TPAMI’23 regression-based one-stage whole body - - 74.2 45.3 87.0 Code
PointHMR [245] CVPR’23 vertex-relevant feature extraction 48.3 32.9 73.9 44.9 85.5 -
PLIKS [246] CVPR’23 inverse kinematics 47.0 34.5 60.5 38.5 73.3 Code
ProPose [247] CVPR’23 learning analytical posterior probability 45.7 29.1 68.3 40.6 79.4 Code
POTTER [248] CVPR’23 pooling attention transformer 56.5 35.1 75.0 44.8 87.4 Code
PoseExaminer [249] ICCV’23 automated testing of out-of-distribution - - 74.5 46.5 88.6 Code
MotionBERT [156] ICCV’23 pretrained human representations 43.1 27.8 68.8 40.6 79.4 Code
3DNBF [200] ICCV’23 analysis-by-synthesis approach - - 88.8 53.3 - Code
FastMETRO [134] ECCV’22 efficient architecture 52.2 33.7 73.5 44.6 84.1 Code
CLIFF [129] ECCV’22 multi-modality inputs 47.1 32.7 69.0 43.0 81.2 Code
PARE [130] ICCV’21 part-driven attention - - 74.5 46.5 88.6 Code
Graphormer [131] ICCV’21 GCNN-reinforced transformer 51.2 34.5 74.7 45.6 87.7 Code
PSVT [133] CVPR’23 spatio-temporal encoder - - 73.1 43.5 84.0 -
@@
GLoT [142] CVPR’23 short-term and long-term temporal correlations 67.0 46.3 80.7 50.6 96.3 Code
MPS-Net [132] CVPR’23 temporally adjacent representations 69.4 47.4 91.6 54.0 109.6 Code
MAED [140] ICCV’21 multi-level attention 56.4 38.7 79.1 45.7 92.6 Code
Lee et al. [161] ICCV’21 uncertainty-aware 58.4 38.4 92.8 52.2 106.1 -
TCMR [139] CVPR’21 temporal consistency 62.3 41.1 95.0 55.8 111.3 -
VIBE [138] CVPR’20 self-attention temporal network 65.6 41.4 82.9 51.9 99.1 Code
ImpHMR [250] CVPR’23 implicitly imagine person in 3D space - - 74.3 45.4 87.1 -
SGRE [163] ICCV’23 sequentially global rotation estimation - - 78.4 49.6 93.3 Code
PMCE [5] ICCV’23 pose and mesh co-evolution network 53.5 37.7 69.5 46.7 84.8 Code
@@
HumanEva dataset [229] is a multi-view 3D human pose estimation dataset comprising two versions: HumanEva-I and HumanEva-II. In HumanEva-I, the dataset includes around 40,000 multi-view video frames captured from seven cameras positioned at the front, left, and right (RGB) and four corners (Mono). Additionally, HumanEva-II features approximately 2,460 frames, recorded with four cameras at each corner.
@@
CMU-Panoptic dataset [230, 251] includes 65 frame sequences, approximately 5.5 hours of footage, and features 1.5 million 3D annotated poses. Recorded via a massively multi-view system equipped with 511 calibrated cameras and 10 RGB-D sensors featuring hardware-based synchronization, this dataset is crucial for developing weakly supervised methods through multi-view geometry. These methods address the occlusion problems commonly encountered in traditional computer vision techniques.
@@
Multiperson Composited 3D Human Pose (MuCo-3DHP) dataset [118] serves as a large-scale, multi-person occluded training set for 3D human pose estimation. Frames in the MuCo-3DHP are generated from the MPI-INF-3DHP dataset through a compositing and augmentation scheme.
@@
SURREAL dataset [231] is a large synthetic human body dataset containing 6 million RGB video frames. It provides a range of accurate annotations, including depth, body parts, optical flow, 2D/3D poses, and surfaces. In the SURREAL dataset, images exhibit variations in texture, view, and pose, and the body models are based on the SMPL parameters, a widely-recognized mesh representation standard.
@@
3DOH50K dataset [153] offers a collection of 51,600 images obtained from six distinct viewpoints in real-world settings, predominantly featuring object occlusions. Each image is annotated with ground truth 2D and 3D poses, SMPL parameters, and a segmentation mask. Utilized for training human estimation and reconstruction models, the 3DOH50K dataset facilitates exceptional performance in occlusion scenarios.
@@
3DCP dataset [232] represents a 3D human mesh dataset, derived from AMASS [233]. It includes 190 self-contact meshes spanning six human subjects (three males and three females), each modeled with an SMPL-X parameterized template.
@@
AMASS dataset [233] constitutes a comprehensive and diverse human motion dataset, encompassing over 11,000 motions from 300 subjects, totaling more than 40 hours. The motion data, accompanied by SMPL parameters for skeleton and mesh representation, is derived from a marker-based MoCap system utilizing 15 optical markers.
@@
DensePose dataset [234] features 50,000 manually annotated real images, comprising 5 million image-to-surface correspondence pairs extracted from the COCO [252] dataset. This dataset proves instrumental for training in dense human pose estimation, as well as in detection and segmentation tasks.
@@
UP-3D dataset [235] is a dedicated 3D human pose and shape estimation dataset featuring extensive annotations in sports scenarios. The UP-3D comprises approximately 8,000 images from the LSP and MPII datasets. Additionally, each image in UP-3D is accompanied by a metadata file indicating the quality (medium or high) of the 3D fit.
@@
THuman dataset [236] constitutes a 3D real-world human mesh dataset. It includes 7,000 RGBD images, each featuring a textured surface mesh obtained using a Kinect camera. Including surface mesh with detailed texture and the aligned SMPL model is anticipated to significantly enhance and stimulate future research in human mesh reconstruction.